### v1

# People who frequently use ChatGPT for writing tasks are accurate and robust detectors of AI-generated text

Abstract
In this paper, we study how well humans can detect text generated by commercial LLMs (GPT-4o, Claude-3.5-Sonnet, o1-Pro). We hire annotators to read 300 non-fiction English articles, label them as either human-written or AI-generated, and provide paragraph-length explanations for their decisions. Our experiments show that annotators who frequently use LLMs for writing tasks excel at detecting AI-generated text, even without any specialized training or feedback. In fact, the majority vote among five such “expert” annotators misclassifies only 1 of 300 articles, significantly outperforming most commercial and open-source detectors we evaluated even in the presence of evasion tactics like paraphrasing and humanization. Qualitative analysis of the experts’ free-form explanations shows that while they rely heavily on specific lexical clues (“AI vocabulary”), they also pick up on more complex phenomena within the text (e.g., formality, originality, clarity) that are challenging to assess for automatic detectors. We release our annotated dataset and code to spur future research into both human and automated detection of AI-generated text.
[…]
[regular] https://github.com/jenna-russell/human_detectors
People who frequently use ChatGPT for writing tasks
are accurate and robust detectors of AI-generated text
Jenna Russell1 Marzena Karpinska2 Mohit Iyyer1,3 1University of Maryland, College Park  2Microsoft  3UMass Amherst {jennarus,miyyer}@umd.edu, mkarpinska@microsoft.com
1 Introduction
Text generated by large language models (LLMs) is rampant in today’s world. [^1] For example, recent studies estimate that over 5% of recently-published Wikipedia articles Brooks et al. (2024) and 10% of PubMed abstracts published in 2024 Kobak et al. (2024) were written by AI. This state of affairs has spurred increased research into automatic detection of AI-generated text, which helps combat academic plagiarism Zhu et al. (2024) and fake content creation Gameiro et al. (2024). Unfortunately, automatic detectors suffer from low detection rates, poor robustness to evasion attempts, and limited explainability to end users Sadasivan et al. (2024); Ji et al. (2024). [^2] This paper focuses only on post-hoc detection of AI-generated text, which unlike watermarking Kirchenbauer et al. (2023) requires no cooperation from LLM providers.
Figure 1: A human expert’s annotations of an article generated by OpenAI’s o1-Pro with humanization. The expert provides a judgment on whether the text is written by a human or AI, a confidence score, and an explanation (including both free-form text and highlighted spans) of their decision.
In this paper, we instead study how well humans can detect AI-generated text. Unlike prior work on this topic, which was mainly conducted in the pre-ChatGPT era Ippolito et al. (2020); Brown et al. (2020); Clark et al. (2021), we focus on text generated by modern LLMs (GPT-4o, Claude-3.5-Sonnet, o1-Pro) and in the presence of evasion attempts (paraphrasing, humanization). We hire human annotators to read non-fiction English articles, label them as written by either a human or by AI, and provide a paragraph-length explanation of their decision-making process (Figure 1). Overall, we collect 1790 annotations on 300 unique articles at a total cost of \$4.9K USD, which allows us to compare humans to automatic detectors and analyze what kind of clues they rely on.
Experience matters:
Unsurprisingly, annotators who rarely or never use LLMs are poor detectors of AI-generated text. However, we identify a subset of five high-performing annotators who frequently use LLMs for writing-related tasks. The majority vote of this subset of “expert” annotators fails to predict the correct label on only one out of 300 articles. We emphasize that we do not provide any training to our annotators: they are given no feedback on which articles they labeled incorrectly.
Human experts outperform automatic detectors:
The majority vote of our five expert annotators substantially outperforms almost every commercial and open-source detector we tested on these 300 articles, with only the commercial Pangram model Emi and Spero (2024) matching their near-perfect detection accuracy. In our most complex configuration, which requires detecting articles generated and humanized by o1-Pro, the expert majority vote is perfect (true positive rate of 100%), while open-source methods such as Binoculars (6.7%) and Fast-DetectGPT (23.3%) struggle. We conclude that hiring expert human annotators to perform detection is a viable strategy, particularly in high-stakes settings where explainability is critical.
What do expert annotators focus on?
An analysis of our expert’s explanations reveals that usage of “AI vocabulary” (e.g., vibrant, crucial, significantly) form the most common giveaways. Close behind are formulaic sentence and document structures (e.g., optimistically vague introductions and conclusions) and originality (how creative or engaging an article is), while complex phenomena such as factuality and tone are also useful signals. We observe that neither paraphrasing nor humanization effectively removes all of these signatures; that said, these evasion tactics and defenses for them are still underexplored in the research community.
Can LLMs mimic human detectors?
We attempt to automatically replicate the decision-making process of our human experts by providing a candidate text to an LLM, along with a guidebook compiled from human expert explanations, and instructing it to use the guidebook to judge whether the text is written by an AI or by a human. This simple prompting strategy is competitive with existing detectors on easier configurations but struggles to detect humanized articles. Moving forward, we believe LLMs can be used in conjunction with high-performing detectors such as Pangram to offer an explanation of the judgment to end users.
Contributions:
Our study establishes the existence of an annotator population (frequent users of LLMs for writing-related tasks) that is highly capable of detecting AI-generated text without any additional training. The majority vote of five such annotators outperforms all automatic detectors we evaluated, even on paraphrased or humanized AI-generated articles. We release our annotated data and codebase to facilitate future research into both human and automatic detection of AI-generated text.
2 How good are humans at detecting AI-generated text?
Automatic detectors of AI-generated text suffer from low detection rates and poor robustness to simple perturbations of the text (e.g., paraphrasing), rendering them unreliable for real-world deployment Sadasivan et al. (2024). How well do human detectors stack up against their automatic counterparts? What features of a candidate text do humans find most predictive of AI writing? To answer these questions, we conducted five experiments that ask humans to judge increasingly more deceptive AI-generated texts (i.e., produced using more powerful base LLMs and evasion tactics).
Task setup:
Each experiment consists of a batch of 60 articles, 30 of which are written by humans and the remaining 30 generated by an LLM. Each human-written article has a corresponding AI-generated counterpart with the same title and subtitle to control for high-level content and topics. The 60 articles are shown to annotators one at a time in a randomized order, and they are asked to provide the following for each article, depicted in Figure 1:
  1. 1.
    a binary label of whether the article was written by a human or an AI
  2. 2.
    a rating of their confidence in their choice on a scale from 1 (least confident) to 5 (most confident)
  3. 3.
    highlighted spans from the article that they used as clues to make their choice
  4. 4.
    a paragraph-length explanation of their choice
Article selection:
We restrict our study to English nonfiction articles of fewer than 1K words geared towards lay audiences. [^3] Human-written articles in our experiments come from eight different American publications: Associated Press, Discover Magazine, National Geographic, New York Times, Reader’s Digest, Scientific American, Smithsonian Magazine, and Wall Street Journal. While we cannot make claims about human detection performance in other domains (e.g., scientific papers or social media posts), [^4] We do report results on §2.1 for fictional stories, which can be found in §C we focus on such articles because (1) malicious LLM users generate large volumes of fake articles, making it practically impactful; (2) such articles do not require specialized knowledge to understand or judge; and (3) many previous studies also ask humans to judge AI-generated text in this domain Ippolito et al. (2020); Brown et al. (2020); Clark et al. (2021); Puccetti et al. (2024).
Generating paired AI articles:
For each human-written article, we generate a corresponding AI article by prompting an LLM with the title, subtitle, desired length, and publication source of the original article. This ensures that AI-generated articles are directly comparable in topic and length, eliminating confounds from content variation. By constructing these human-AI article pairs with only authorship as the varying factor, we create minimal pairs that allow direct comparisons Gardner et al. (2020); Warstadt et al. (2020); Karpinska et al. (2022, 2024). We employ a within-subjects design for our experiments, where each annotator judges both the human-written and AI-generated articles. This reduces variability from individual differences and requires fewer annotators Allen (2017). To further minimize bias, annotators are unaware of the pairing, and the article order is randomized. An example article generation prompt is below:
You are given the following title and subtitle of a general article from the Science section of New York Times and asked to write a corresponding article of around 750 words. Include quotations from relevant experts and make sure the article is concise and easily understandable to a lay audience.
Title: The Science That Makes Baseball Mud ‘Magical’
Subtitle: Scientists dug up the real dirt on the substance applied to all the baseballs used in the major leagues.
Article:
For experiments 3 and 5, we modify the prompt to include perturbations designed to evade detectors (paraphrasing and humanization). Comparisons of article lengths between human and AI-generated articles for each experiment are reported in Table 1 and further detailed in §B.1.555We set temperature=0 for Experiments 1-3 and use ChatGPT (Pro) interface for Experiments 4-5 as the API for o1 was not yet available.
Exp # Model Human-written (n=30) AI-generated (n=30)
1 GPT-4o 652.5170.9 625.1126.9
2 Claude-3.5-Sonnet 793.1170.3 710.8133.5
3 GPT-4o paraphrased 777.8175.9 654.9110.6
4 o1-Pro 728.4155.1 813.8233.6
5 o1-Pro humanized 739.0128.3 815.6171.2
Table 1: Mean and standard deviation (subcripted) of article length in words across experiments, computed by splitting on whitespaces.
Annotator details:
We recruited annotators via Upwork, [^6] https://www.upwork.com/ all of whom identified themselves as native English speakers. All annotators were required to read the project guidelines and sign a consent form prior to the task, provided in §A. Additionally, we surveyed each annotator to determine their level of education, profession, English dialect, and familiarity with LLMs, including how they use these models (see Table 5 for details). They were compensated \$2 per candidate text, resulting in an hourly rate of about \$15 - \$30 USD. Ultimately, we collect 1740 annotations from 9 annotators at a total cost of \$4.9K USD.
Evaluation metrics:
We evaluate both human and automatic detectors on True Positive Rate (TPR), which measures the percentage of AI-generated articles that are successfully detected, as well as False Positive Rate (FPR), which measures the percentage of human-written articles incorrectly detected as AI-generated. Prior work in automatic detection reports either TPR at a fixed low FPR such as 1% Krishna et al. (2023); Hans et al. (2024); Dugan et al. (2024) or AUROC Mitchell et al. (2023); Hu et al. (2023); Bao et al. (2023), which we cannot easily do with human annotators. For fair comparison, we report TPR and FPR of automatic detectors using recommended thresholds whenever possible, otherwise using a threshold calibrated to an FPR of 5% on a held-out set of 40 human-written articles.
[…]
Is there a population of humans that is highly accurate at our detection task, and if so, are there any commonalities between them? To explore these questions, we first ask five annotators with varying backgrounds and levels of expertise with LLMs to attempt a batch of 60 articles (half generated by GPT-4o 2024-08-06 and half human-written as described above). We observe both ends of the spectrum in terms of performance: annotators who rarely or never use LLMs for writing tasks performed near random, while one annotator who uses LLMs (and edits LLM-generated text) daily performed almost perfectly. More details on our initial study can be found in §C. Inspired by this result, we recruit four more annotators with similar backgrounds as our top performer, all of whom perform similarly well on the same batch of articles.
Annotators with limited LLM experience are unreliable detectors:
The four annotators who self-report either not using LLMs at all, or using LLMs only for non-writing tasks, detect AI-generated text at a similar rate to random chance, achieving an average TPR of 56.7% and FPR of 52.5% (Table 2). [^7] We report the average TPR and FPR since there were only four non-expert annotators. For the remainder of the paper, we will report the majority vote of the five expert annotators (i.e., at least three out of five agreed). While performance among these annotators varies greatly, no single annotator could reliably detect AI-generated text. We term this population nonexperts for the rest of this paper.
Annotators who frequently use LLMs for writing tasks are highly accurate:
In contrast, the five annotators who have significant experience with using LLMs for writing-related tasks are able to detect AI-generated text very reliably, achieving a TPR of 92.7%. The average FPR for this population of annotators was 3.3%, meaning that they rarely mistake human-written text as AI-generated. The majority vote out of these five annotators correctly determined authorship of all 60 articles; the “GPT-4o” column of Table 3 contains more details. We term this population experts for the rest of this paper; all subsequent experiments rely on only these five expert annotators.
What do expert annotators see that nonexperts don’t?
To understand why experts far outperformed nonexperts at detecting AI-generated text, we analyze the comments each annotator provided in their explanations. Overall, nonexperts often mistakenly fixate on certain linguistic properties compared to experts. One example is vocabulary choice, where nonexperts take the inclusion of any “fancy” or otherwise low-frequency word types as signs of AI-generated text; in contrast, experts are much more familiar with exact words and phrases overused by AI (e.g., testament, crucial). [^8] A complete list of “AI vocab” found in the detection guide (Table 11) is detailed in Table 12. Nonexperts also believe that human authors are more likely to form grammatically-correct sentences and thus attribute run-on sentences to AI, while experts realize the opposite is true: humans are more likely than AI to use ungrammatical or run-on sentences. Finally, nonexperts attribute any text written in a neutral tone to AI, which results in many false positives because formal human writing is also often neutral in tone.
[…]
Avg. TPR 56.7 92.7
Avg. FPR 51.7 4.0
Avg. Confidence 4.027 4.394
Table 2: On average, nonexperts perform similar to random chance at detecting AI-generated text, while experts are highly accurate.
[…]
$2 \faFlask$ Experiment 2: Can experts detect articles generated by a different LLM?
[…]
While ChatGPT is the most widely-used LLM (both in general and by our expert annotators), competitors such as Anthropic’s Claude Anthropic (2023) also have a wide user base. Are our experts overfitting to artifacts specific to GPT-4o, or do they detect patterns that generalize to other LLMs? We address this question in Experiment 2 by evaluating our experts on a second batch of 60 articles, this time with 30 new human-written articles and 30 corresponding articles generated via Claude-3.5-Sonnet Anthropic (2024) using the same instructions as in Experiment 1.
Experts reliably detect articles generated by Claude:
Despite the change in model, experts are reliable at detecting AI-generated content. TPR is almost unchanged from Experiment 1 and the expert majority vote is again 100% accurate (“Claude-3.5-Sonnet” column of Table 3), with two annotators achieving a perfect score individually. We note that one expert (Annotator 2) achieved higher performance in the experiments where an OpenAI model was used, with their TPR falling to 80% in this experiment. Annotator 2 did not have any prior experience using Claude, although three other annotators who were also unfamiliar with Claude achieved TPRs of at least 96.7%. However, Annotator 2 focuses more heavily on the existence of “AI vocab” in the article than the other annotators, and they were not familiar with the distribution of words that are overused by Claude.
[…]
Users of LLMs who hope to evade detection often resort to simple post-hoc modifications of LLM outputs such as paraphrasing. These tactics significantly drop detection rates of automatic methods Krishna et al. (2023); Sadasivan et al. (2024), but how do they affect our human experts? We evaluate the five expert annotators on a third set of 60 articles, where the 30 AI-generated articles go through an additional round of paraphrasing, and we show that they maintain high TPR and low FPR.
Creating paraphrased AI-generated articles:
We follow the sentence-by-sentence paraphrasing approach with LLMs outlined by Chang et al. (2024) to create a more difficult set of AI articles for this experiment. Additionally, since our experts easily identified AI-generated named entities (e.g., Sarah Thompson) in the previous experiments, we extract named entities from the human-written article and instruct GPT-4o to include these names in its generated article. We use GPT-4o as both the article generation model and the paraphrasing model. See Table 10 for paraphrasing prompts used in this experiment.
Experts are robust to paraphrased AI-generated articles:
Overall, TPR and FPR remain almost unchanged in this experiment compared to the first two, suggesting that paraphrasing is not an effective attack on expert human detectors (see “gpt-4o paraphrased” column in  Table 3). The majority vote among the five experts again correctly detects the labels of all 60 articles. Annotators continue to pick up on many of the same clues within the paraphrased articles that were also apparent in Experiments 1 & 2, such as high frequency of “AI vocab” (even after paraphrasing), formulaic sentence structures, and cheerful summary conclusions.
[…]
While we were in the middle of conducting experiments, OpenAI released their o1 model OpenAI (2024), ushering in a new paradigm of test-time scaling. This release offered us a unique opportunity: since none of our experts had been previously exposed to a model with such advanced reasoning capabilities Zhong et al. (2024), how well would they do at detecting its output? Interestingly, our experts remain reliable detectors of articles generated by o1-Pro (using the same prompt as in Experiments 1 and 2),999o1-Pro is available through the ChatGPT Pro subscription. OpenAI states o1-Pro produces more reliably accurate and comprehensive responses than o1. although their comments show that they often perform a more detailed analysis of articles to make their judgments.
Four out of five experts are robust to o1-Pro:
For four of five experts, detection ability of o1-generated content remains largely consistent with prior experiments, with a majority vote TPR of 96.7% and FPR of 0% (see “o1-Pro” column of Table 3). [^10] One AI-generated article was incorrectly marked by the expert majority vote as human-written. Analyzing their explanations reveals that the three annotators who rely most heavily on vocabulary and sentence structure did not find enough clues in those categories to judge the article as AI-generated. The full text of the article is in Table 14. Interestingly, Annotator 3’s TPR drops considerably, as they prioritize signs of human writing over signs of AI writing when making their judgments. Average confidence dropped to 4.21 out of 5, compared to average confidence of 4.39, 4.38, and 4.48 from Experiments 1,2, & 3 respectively (see Figure 2 for details). The drop in confidence, as well as verbal feedback from our annotators, demonstrates the increased challenge posed by o1-generated articles. However, their steady aggregate performance highlights that even new model paradigms are still detectable by human experts.
Experts provide more nuanced explanations when detecting o1-Pro:
In prior experiments, expert explanations focused primarily on whether or not a candidate text possesses characteristics of AI. However, experts shift focus with o1-Pro by more frequently commenting on identifying characteristics that make text sound “human”. For instance, experts frequently point to how humans repeatedly write the word says, while AI tries to use more descriptive synonyms like notes and explains. Annotator 3, the only one consistently fooled by o1-Pro outputs, relied too much on signs of informality (e.g., contractions, slang usage, usage of just and actually) as a sign of human writing, with 66.7% of their explanations relating to formality (Figure 11).
[…]
As the deployment of automatic AI detectors increases, many users are attempting to evade them by using humanization methods, which explicitly attempt to make AI-generated text more human-like Wang et al. (2024a, b) unlike more generic attacks such as paraphrasing. Are such methods effective at evading our human experts? Since no well-established humanization methods exist, we first create our own humanizer by prompting o1-Pro with a detailed set of instructions derived from expert comments written for Experiments 1-4. Despite considerably degrading performance of many automatic detectors, the majority of our experts remain robust to humanization (“o1-Pro humanized” column of Table 3).
Implementing a prompt-based humanizer:
We modify the article generation prompt used in prior experiments to include detailed instructions on specific AI signatures to avoid when generating the article. To obtain these instructions, we consult our experts directly, paying them \$45 each to provide us with a list of clues that they look for during detection. We then manually organize these disparate clues into a unified “guidebook” with different sections (e.g., vocabulary, grammar, tone, introductions, conclusions), where each section provides explanations and examples of how AI writing differs from human writing (see Table 11 for a truncated version of the guidebook). Then, we prompt o1-Pro with the guidebook and an instruction to generate an article that would evade a detector that was following the guidebook (see Table 13 for prompt). [^11] We only use o1-Pro for this experiment as the previous experiments show it was more difficult to detect than GPT-4o and Claude-3.5-Sonnet. Through this process, we create a new batch of 60 articles, 30 of which are generated by our humanization prompt using o1-Pro. More details on the development of our humanization method can be found in §B.2.
Experts remain robust to humanized articles:
Despite our best efforts to generate articles that our experts would find undetectable, most of their detection rates remain largely unchanged from prior experiments, and the expert majority vote is again perfect on all 60 articles. Annotator 3, who struggled to detect non-humanized o1-Pro articles in Experiment 4, performed remarkably poorly on this batch, achieving a TPR of zero by marking almost every article as human-written. While detection rates remained steady, overall confidence dropped, with 15.1% of annotations having a confidence of 1, showing the increased difficulty presented by the humanized articles (see Figure 2).
Figure 2: Expert confidence in their decisions drops when judging humanized articles generated by o1-Pro.
AI clues still remain after humanization:
Although our experts reported this round to be the most challenging, they remained consistent in the types of clues used from prior rounds, demonstrating that the humanization prompt did not remove the entirety of the “AI signature” from the articles. For example, while humanization increases variety in AI-generated names, often even including people relevant to the article who exist in real life, our experts found that humanized articles use titles for these people (e.g., Dr., Prof.) much more frequently than in human-written articles. Overall, our results highlight the need for increased public research on humanization, as more advanced methods can challenge both human and automated detectors while offering better training data for more robust detection systems.
3 Fine-grained analysis of expert performance
Category Freq Definition Example Explanations
Vocabulary 53.1% LLMs use specific words and phrases more often than human writers, which often results in repetitive, unnatural, or overly complex wording. Human: "Furthermore, I very much doubt AI would have used adventurous adjectives like ‘chunky’, ‘musky’ or ‘thin’ to describe food. Nor would it have used verbs like ‘blitzing’ or ‘bolstering’."
AI (o1-Humanized): "Odd word choices: wheat that ‘stores’ a lineage; genes that are ‘honed.’"
Sentence Structure 35.9% AI-generated sentences follow predictable patterns (e.g., high frequency of “not only … but also …”, or consistently listing three items), while human-written sentences vary more in terms of length. Human: "Short choppy sentences and paragraphs."
AI (o1-Pro): "One pattern I’ve been noticing with AI, and I think I’ve stated this before, is the comparison of ‘it’s not just this, it’s this’ and I’m seeing it here, along with listings of specifically three ideas."
Grammar & Punctuation 24.8% AI-generated text is usually grammatically perfect (also avoiding dashes and ellipses), while human-written text often contains minor errors. Human: "There’s a lot of variety in the article’s grammar use, with dashes, brackets, quotes intermixed with sentences, and short spurts of comma sections throughout."
AI (GPT-4o-Para):"there’s nothing off about the grammar or syntax in this piece…"
Originality 23.7% AI-generated writing is generally straightforward, “safe,” and lacking in surprises or humor, leaving annotators bored or disengaged. Human: "it’s offset by some great analogies and creative phrasing that works well to convey the topic, such as with "amateur sleuths", "catnip for a certain type of Reddit user."
AI (o1-Pro): "What happens when AI tries to be creative? Penguins "stand on their own flippers"."
Quotes 22.3% AI-generated quotes sound overly formal, lack the varied nuances of real conversation, and often mirror the article’s main text too closely in style. Human: " The quotes being short snippets also makes me think they’re real, as the writer had to find a way to fit them into the text, rather than them just perfectly stating either side’s views."
AI (GPT-4o): "The quotes also feel fake, every expert speaks the same way and it’s too homogenous with the text."
Clarity 19.5% AI-generated text often lacks concise flow by over-explaining or including irrelevant details, effectively “telling” rather than “showing”. Human: "Words like "meander" are used, but are used sparingly to create better flow of ideas, and its writing style is simplified in the best way possible."
AI (Claude-3.5-Sonnet): "The sentences are condensed to provide the best possible precision with its word choice, but the article lacks flow and clarity."
Formatting 15.0% AI-generated formatting is overly consistent (e.g., fully capitalized headings, bolded lists, paragraphs of similar length). Human: ""the formatting itself is varied, with topic headers focusing on lowercase lettering, longer explanations, and the bullet points not going into a title: one to two sentence explanation format … "
AI (o1-Pro): "… it has a structured article format of commonly used headers…"
Conclusions 13.1% AI-generated conclusions tend to be repetitive and overly optimistic summaries, while human-written conclusions end more abruptly and less tidily. Human: "Both the introduction and the conclusion were humorous and unique."
AI (Claude-3.5-Sonnet): Most of the conclusion leads to that summarizing, flowery tone…
Formality 12.3% AI-generated text (especially without humanization) rarely contains filler words, contractions, slang, or abbreviations, favoring fully spelled-out terms and a polished tone. Human: "It includes filler words like ‘just’, ‘very’ and ‘really’. Colloquial language like ‘sucked’."
AI (Claude-3.5-Sonnet): "Phrases like ‘maintains that extending habeas corpus rights to animals’ and ‘fundamentally alter the ability of accredited zoos to conduct their vital conservation work.’ are so strictly formal, so tight and dense with wordage."
Names & Titles 11.7% LLMs frequently generate the same names regardless of the prompt (e.g., “Emily Carter,” “Sarah Thompson”). Furthermore, if an article contains multiple people, they tend to refer to all of them with the same title (e.g., “Dr.”) without offering unique details. Unique or context-specific names (including real brands or products) are associated with human writing. Human: "A couple of the experts also have quite unique names, and none of them are referred to as Dr. X (bonus points for none of them being named Emily)."
AI (GPT-4o): "However, the introduction of Dr. Sarah Thompson and Dr. Emily Carter (who by now has more than a lifetime’s worth of qualifications), means that it has to be AI text."
Tone 9.3% The tone of AI-generated text is consistently neutral or positive, lacking depth or emotional variety compared to human writing. Human: "The article seems to display some bias against TikTok while trying to remain impartial or impartial-seeming. The same goes for its political slant which although subtle seems to be there."
AI (o1-human): "…the majority of the writing is filled with the same language it uses to describe everything - inspirational, stunning, essential, and resonating - using formal words, an inherent positivity bias, and a reflective, romantic tone that doesn’t give details on why this topic matters, why we still don’t know who she is, and whether or not there was any kind of controversy around the idea of women being on stage…"
Introductions 7.3% AI-generated introductions are usually generic or focus on scenic details without providing key background, while human introductions have more compelling hooks and relevant context. Human: "The introduction is unique as it starts with the subject of the article watching a movie rather than instantly explaining the entire point of the article."
AI (Claude-3.5-Sonnet): "The article has a very generic introduction and conclusion, especially the introduction, which essentially tells you what the entire article is about."
Factuality 7.2% AI-generated text contains factual inconsistencies or hallucinations more often than human writing. Human: "The article is very factual, with details about the dates, materials used, and sales portions."
AI (GPT-4o-Para): "Incorrect information delivered with customary AI confidence: the ‘quad axel’ is not another name for the backflip."
Topics 3.1% Unlike human authors, LLMs generally avoid darker or more mature topics (e.g., violence, graphic descriptions). Human: "Phrases throughout the article, including "burned to ashes and scattered at sea to prevent the crowd from venerating them as relics" and "to trample on a brass likeness of Jesus or the Virgin Mary—a blasphemous act." actively show what happened, the horrors and tragedies that occurred during that time."
AI (GPT-4o): "It spends very little time talking about the horrors of the disease, and instead focuses on future research, hopeful quotes, and potential cures, even referring to it as "embarking on a new chapter"."
Other 2.6% Other clues that do not fall into any of the above categories, often based on the annotator’s intuition or overall impression. Human: "There are no clues here apart from the highlighted sentence which seems to have a human ‘ring’ to it."
AI (o1-Humanized): "The article feels slightly artificial but I cant́ really find any clear clues for it."
Table 4: Taxonomy of clues used by experts to explain their detection decisions. For each category, we report the frequency of explanations that mention that category (regardless of if the annotator was correct) and provide examples of explanations for both human-written and AI-generated articles. While vocabulary and sentence structure form the most frequent clues, more complex phenomena like originality, clarity, formality, and factuality are also distinguishing features.
The previous section establishes that human experts are highly accurate detectors of AI-generated text even in the presence of evasion attempts. Here, we first provide in-depth comparisons between experts and automatic detectors. Next, we perform a fine-grained coding of expert explanations into different categories (e.g., vocabulary, originality, formality), which allows us to examine what they focus on during the different experiments. Finally, we analyze differences between annotators and what they focus on when they are incorrect, highlighting implications for the future training of human annotators for AI-generated text detection.
Human experts vs. automatic detectors:
We begin with a direct comparison between our human experts and state-of-the-art AI detectors. Specifically, we run five different detectors on our set of 300 articles:
  • Pangram Emi and Spero (2024) is a closed-source commercial detector implemented via a Transformer classifier trained via an iterative process that uses hard negative mining and synthetic data to improve data efficiency. We run both their base model and a newer model (Pangram Humanizers) trained to distinguish humanized data Masrour et al. (2025). We directly use the labels (i.e., thresholds) produced by the Pangram API and do not award credit for the neutral label “Possibly AI”.
  • GPTZero Tian and Cui (2023) is also a closed-source commercial detector that runs a classifier sentence-by-sentence across the document. We directly use the binary labels produced by the GPTZero API.
  • Binoculars Hans et al. (2024) is an open-source detector that relies on the cross perplexity computed by two different language models to perform detection. We run Binoculars with the two FPR thresholds recommended by its authors (“Accuracy” and “Low FPR” modes).
  • Fast-DetectGPT Bao et al. (2023) is an open-source method that samples and scores many perturbations of the text to estimate conditional probability curvature. We threshold this method at a FPR of 0.05, set on our held-out dev set of 40 articles.
  • RADAR Hu et al. (2023) train an open-source classifier adversarially against a paraphraser. We threshold this method at a FPR of 0.05, set on our held-out dev set of 40 articles.
The middle rows of Table 3 show that only Pangram Humanizers (average TPR of 99.3% with FPR of 2.7% for base model) matches the human expert majority vote, and it also outperforms each expert individually. Pangram is near perfect on the first four experiments and falters just slightly on humanized o1-Pro articles, while GPTZero struggles significantly on o1-Pro with and without humanization. The open-source detectors degrade in the presence of paraphrasing and underperform both closed detectors by large margins on average. We again note that our experts are untrained at this detection task, and they could likely improve their individual performance if provided with feedback on their errors.
Coding expert explanations:
Our automatic baselines are largely unexplainable: they either rely on combinations of various statistical properties of the text, or operate as opaque classifiers trained on large datasets. In contrast, we can easily solicit explanations from human annotators, and in this section we use GPT-4o to code these explanations into a schema (Table 4) developed by the authors after careful manual analysis. Details of the coding process used to label explanations can be found in §D.
Vocabulary-related clues are mentioned in the majority (53.1%) of all explanations, while sentence structure (35.9%), grammar (24.8%), and originality (23.7%) are also common. We note that many of these categories (e.g., originality, factuality, tone) are much more difficult to assess automatically than others (e.g., vocabulary), and these may currently be areas where humans have an advantage over automatic detectors.
Figure 3: (Top) A heatmap displaying the frequency with which annotators mentioned specific categories in their explanations when they were correct. Interestingly, vocabulary becomes a less frequent clue for o1-Pro-generated articles, especially with humanization. (Bottom) Same as above, except only computed over explanations when experts were incorrect. Formality is a big source of misdirection for o1-Pro articles, while fixating on sentence structure can lead experts to false positives. Details of each category can be found in Table 4.
When experts correctly detect AI-generated articles, what clues do they use?
Figure 3 (upper) shows the frequency that clue categories are mentioned in explanations for which the expert makes the correct decision. We observe several interesting shifts as the complexity of the article generation setup increases: for example, 57.1% of correct explanations about o1-Pro articles mention vocabulary, compared to only 42.3% for the humanized o1-Pro articles prompted to avoid “AI vocab”. Somewhat counterintuitively, explanations about paraphrased articles note AI vocab in 88% of explanations, compared to only 69.8% of non-paraphrased GPT-4o articles. Similarly, quotations are mentioned in 33.8% of explanations about paraphrased articles, a much higher rate than other configurations: a close reading of explanations reveals that experts flagged quotes that were always in the same format and style (e.g., only placed at the end of each paragraph). Originality is a common clue for detecting human-written text, highlighting the gulf in creativity between humans and LLMs Chakrabarty et al. (2024).
Annotators don’t always focus on the same clues:
Our experts focus on different properties of the text to arrive at their decisions. Annotator 1 is the only one to pick up on “AI names”: in fact, 63.3% of GPT-4o and 70% of Claude-3.5-Sonnet articles include either the name Emily or Sarah.121212o1-Pro favors names of real people instead of fictional names when generating articles. Annotator 2 and 3 emphasize linguistic features, such as humans not always following “proper” writing conventions or the tendency of LLMs to always list examples in groupings of three. Annotator 4 focuses more on the flow of the articles, analyzing the specificity of detail and motivation behind an article, while Annotator 5 frequently mentions whether a quotation sounds natural or not. Further analysis of individual performance can be found in §D.2. This diversity partly explains why an ensemble of expert annotators performs so well, and it also suggests room for training so that all experts are at least aware of most distinguishing features.
When experts are incorrect, what clues lead them astray?
Figure 3 (lower) shows the frequency that clue categories are mentioned in explanations for which experts make incorrect decisions. Despite near-perfect majority vote performance, experts make several errors individually, particularly with the qualitative change in text produced by o1-Pro. For Annotator 3, who relies heavily on linguistic features, o1-Pro’s frequent use of contractions and colloquialisms dramatically lowered their detection rate: they mention formality in none of their explanations for Experiments 2 and 3, compared to 66.7% of o1-Pro and 83.3% of humanized o1-Pro explanations as shown in Figure 11. Studying expert false positives is also insightful Figure 3 (lower): 31% of explanations for false positives mention vocabulary, typically when human-written content contains words like “delve” and “crucial” that our experts associate with AI-generated text. More prominently, 50% of false positives focus on sentence structure, when humans write stylistically-similar sentences to those preferred by LLMs (e.g., contiguous blocks of sentences with similar lengths, multiple lists of three items).
4 Can LLMs be prompted to mimic human expert detectors?
Our experiments so far highlight the benefits of hiring humans to perform AI-generated text detection: in aggregate, they are more accurate and robust than all automatic detectors we tried. Additionally, they can provide detailed explanations of their decision-making process, unlike all of the automatic detectors in our study. However, an obvious drawback is that hiring humans is expensive and slow: on average, we paid \$2.82 per article including bonuses, and we gave annotators roughly a week to complete a batch of 60 articles. In this section, we prompt LLMs to imitate our expert annotators, providing the guidebook used in our humanization experiments (§2.5) to the model and asking it to produce an explanation and a label. While the approach shows promise, outperforming detectors such as Binoculars and RADAR, it fails to reach the performance of our human experts as well as advanced automatic detectors like Pangram.
Implementing a prompt-based detector:
Using our guidebook for AI detection (Table 11), we prompt an off-the-shelf LLM to decide whether a candidate text is human-written or AI-generated and explain its decision based on the criteria in the guidebook (see Table 21 for prompt template). While we can control this detector’s behavior by modifying the prompt (e.g., provide explanations before producing a final label), it does not output a single score, which makes setting a false positive threshold impossible. [^13] We do not experiment with thresholding the token-level probabilities of the labels “AI” vs. “Human”, since these are inaccessible with closed LLMs, but this is a promising route for future research on open prompt-based detectors. We derive the final output of our detector by using the prediction from the detection model in a deterministic configuration.
Prompt-based detection shows promise but struggles with humanization and high false positive rates:
We implement our prompt-based detector by zero-shot prompting GPT-4o-2024-11-20 and o1, [^14] We set temperature=0 for GPT-4o and reasoning efforts to ‘medium’ for o1 (note that OpenAI has fixed temperature=1 for o1). Further experiments can be found in Table 22. comparing the effects of chain-of-thought (CoT) prompting Wei et al. (2022) and the inclusion of the detection guide on its performance. For GPT-4o, the best detector configuration is achieved with both CoT and the guidebook (average TPR 78%), while o1 is best with CoT but without the guidebook (54%). We observe that o1 is more conservative than GPT-4o, demonstrating lower FPRs on average at the cost of lower TPR. However, it is also more capable at detecting humanized text. On average, our best configuration with GPT-4o-2024-11-20 performs comparably to Binoculars and Fast-DetectGPT, although obviously at a much higher cost. The gap between this method and humans demonstrates that we have a long way to go in order to “teach” LLMs how to detect AI-generated text like humans do; we speculate that fine-tuning may be helpful to close the gap.
5 Related work
Our work builds on prior research centering around AI-generated text detection using both human and automatic approaches, as well as recent work on evading such detectors.
Human detection of AI-generated text:
Our study most closely resembles several papers published prior to ChatGPT’s release. Existing work notes that while naïve annotators do not reliably detect AI-generated texts Ippolito et al. (2020); Brown et al. (2020); Clark et al. (2021); Karpinska et al. (2021), some annotators manage to perform very well Ippolito et al. (2020). More recently, these findings have been reproduced in other domains such as Italian news articles Puccetti et al. (2024) and poetry Porter and Machery (2024). Dugan et al. (2020) and Dugan et al. (2022) address a more complex task where humans and machines are tasked to identify boundary between human and AI-generated text in mixed documents.
Automatic detection of AI-generated text:
Successful automatic detection methods are typically either perplexity-based Mitchell et al. (2023); Bao et al. (2023); Hans et al. (2024) or trained classifiers Solaiman et al. (2019); Emi and Spero (2024); Verma et al. (2023). Some detectors have focused on sentence-level detection Kushnareva et al. (2024); Wang et al. (2023), emphasizing the need for more detailed and explainable detection. Dugan et al. (2024), Li et al. (2024), and Zhang et al. (2024) attempt to compare the performance of detectors in different domains and adversarial attack settings. However, most automatic detectors are unreliable in the face of attacks such as paraphrasing Krishna et al. (2023); Sadasivan et al. (2024) and highly sensitive to stylistic factors Doughman et al. (2025).
Evading detection:
Increasing LLM usage among the general population has spurred corresponding interest in humanization of AI-generated text. [^15] Relevant discussions on Reddit: ”AI Humanizer Recommendations?” (2024), ”How to humanize ai-generated texts?” (2024), ”How to humanize the AI generated content?” (2024) In fact, many commercial humanization services exist to fill this demand. [^16] Some humanization services include Undetectable AI, MyEssayWriter.ai and Stealth Writer. Initial research has proposed several humanization attacks to evade automatic detection Wang et al. (2024b); Lu et al. (2023); Shi et al. (2023); Wang et al. (2024a), while Zhou et al. (2024) and Masrour et al. (2025) improve the robustness of automatic detectors to humanized content.
Analyzing differences between human-written and AI-generated text:
Many frameworks group errors in machine-generated text into categories similar to the ones we use in this paper Gehrmann et al. (2019); Dou et al. (2021). More recently, Ma et al. (2023) discover gaps in terms of depth and content quality between scientific AI-generated and human-written text. Shaib et al. (2024) introduces syntactic templates, noting the repetitiveness of AI-generated text in comparison to human-written references. More closely related to our work is that of Ji et al. (2024), who qualitatively categorize human detection explanations but do not employ expert annotators. Jakesch et al. (2023) explores the heuristics humans use to detect AI-generated text, with some similar conclusions to ours (e.g., humans are likely to associate more informal text with human writing).
6 Conclusion
Our paper demonstrates that a population of “expert” annotators—those who frequently use LLMs for writing-related tasks—are highly accurate and robust detectors of AI-generated text without any additional training. The majority vote of five such experts performs near perfectly on a dataset of 300 articles, outperforming all automatic detectors except the commercial Pangram model (which the experts match). Analysis of explanations provided by our expert annotators reveals that they pick up on not just vocabulary and sentence structure-related clues but also more complex properties like originality, factuality, and tone. We observe that each expert organically focuses on different aspects of the text, and we conjecture that with explicit training, human annotators can be made even more robust to advances in LLMs as well as evasion tactics (e.g., humanization). Future work can also explore human annotators working alongside automatic detectors like Pangram to improve detection accuracy and explainability. We find it apt to end with a comment from one of our experts about a particularly formulaic conclusion generated by GPT-4o:
This time I went to the end of the piece and said: “Hello, AI.” There it was in all its glory: the “testament” serving “as a beacon of hope and inspiration” and “demonstrating” to us humans “that anything is possible.” Sometimes I feel sorry for AI—it must have a dreary time trying to satisfy its human interlocutor’s desire to “showcase” advocacy, social change, inclusivity, gender equality, equity, and representation in an essay of fewer than a thousand words. Annotator 5
Limitations
Our study is limited to articles in American English, chosen for their consistent formatting and high quality (i.e., professionally written and proofread). We also did not investigate factual accuracy, as it did not appear to be a significant cue for our annotators, who covered a broad range of topics. Finally, while we selected articles from reputable sources, there remains a possibility that some included AI-generated edits beyond our scope of detection.
Ethical Considerations
This study was reviewed by the UMass Institutional Review Board (IRB #5927) and deemed exempt. All annotators were briefed on the purpose of the project and provided informed consent prior to participating. Those who wished to be acknowledged by name explicitly agreed to do so in their consent forms. We ensured fair compensation for annotators in recognition of their time and expertise. We acknowledge the potential risks of misinformation and hallucinated content, especially when AI outputs are presented as human-written. Our goal is to examine these issues and inform best practices, rather than endorse or facilitate deceptive uses of AI.
Acknowledgments
We would like to extend our gratitude to our Upwork expert annotators for their dedicated efforts, and link their Upwork accounts for future work requiring human expertise on AI text detection: Benjamin Stewart, [^17] Benjamin’s Upwork: https://www.upwork.com/freelancers/~01cfc42298090772b8?mp_source=share Elizabeth Bergman, [^18] Elizabeth’s Upwork: https://www.upwork.com/freelancers/~01e0d3569ef95d391d?mp_source=share Lynette Stewart, [^19] Lynette’s Upwork: https://www.upwork.com/freelancers/~01cc5594cf2405ea5e?mp_source=share Roberta King, [^20] Roberta’s Upwork: https://www.upwork.com/fl/robertaking?mp_source=share and Shannon West. [^21] Shannon’s Upwork: https://www.upwork.com/freelancers/~01daaa89e049285294?mp_source=share We thank Pangram and GPTZero for providing credits to access their APIs. We are also grateful to the members of the UMass NLP lab for their insightful input. This project was partially supported by awards IIS-2202506 and IIS-2312949 from the National Science Foundation (NSF) as well as an award from Open Philanthropy.
Appendix A Human Evaluation
In this section of the appendix we provide additional details about our experts the data collection pipeline.
Annotators:
The annotations for the experiment 1 were done by 5 annotators recruited on Upwork. All annotators are native English speakers from the US or South Africa. All annotators hold university degrees, worked in varying professions, and had varying levels of familiarity with AI assistants like ChatGPT. One had never used AI, 2 had little experience, and 2 used AI every day. Our five expert annotators are native English speakers hailing from the US, UK, and South Africa. Most work as editors, writers, and proofreaders and have extensively used AI assistants. See Table 5 for more information about annotators.
Annotator Education Level English Dialect LLM Usage AI Models Used Occupation
Non-Expert
- Professional Degree or Doctorate American English Few times ChatGPT Writer
- Bachelor’s Degree American English Daily ChatGPT, Gemini, Llama Finance/Operations
- Some college, no degree American English Few times None Transcription, editing, data entry
- Bachelor’s Degree American English Never None English tutor, copywriter, author
Expert
Annotator 1 Bachelor’s Degree South African English Daily ChatGPT, Microsoft Copilot Freelance editing, writing, proofreading
Annotator 2 Bachelor’s Degree South African English Weekly ChatGPT Editing and Proofreading
Annotator 3 Master’s Degree British English Daily ChatGPT Copyeditor and Proofreader
Annotator 4 Bachelor’s Degree American English Weekly ChatGPT, Llama, Huggingface models Freelancer Content Writer
Annotator 5 Master’s Degree South African English Weekly ChatGPT, Claude Language teacher
Table 5: Survey of Annotators, specifically their backgrounds relating to LLM usage and field of work. Note that expert Annotator #1 was one of the original 5 annotators (along with the 4 non-experts) and remained an annotator for all expert trials.
Collecting Human Annotations
All annotators were required to read the guidelines (Figure 4) and sign a consent form (Figure 5) prior to the labeling task. Collecting all labels usually required additional communication with the annotators, resulting in about 20 hours of work from the author involved in this process. We estimate that the annotators were able to read and label between 8 and 12 articles per hour based on self-reported time and records in the spreadsheets. Figure 6 shows the interface annotators use to complete the annotation process. They read and highlight an article, then complete the annotation by providing their decision, confidence score, and explanation. Since completing annotations for 60 article batches takes a long time (estimated 6-8 hours of work), we have implemented an interface that made it possible for the annotators to save their work and come back at any time, allowing annotators to allocate their time as they saw best. Annotators were given 1 week to complete batches of 60 articles, with more time allowed when needed. Note that unlike Figure 1, annotators did not see the titles of the article when completing annotations. This decision was made to prevent obvious linkage of AI-generated and human-written text pairs.
(a) Guidelines: Page 1
(b) Guidelines: Page 2
(c) Guidelines: Page 3
(d) Guidelines: Page 4
Figure 4: Guidelines provided to the annotators for the annotation task. The annotators were also provided additional examples and guidance during the data collection process.
Figure 5: Consent form which the annotators were asked to sign via GoogleForms before collecting the data.
Figure 6: Interface for annotators, with an example annotation from Annotator #4 with a humanized article from §2.5. This is the same article displayed in Figure 1. An annotator can highlight texts, make their decision, put confidence, and write an explanation. This AI-generated article was based off of In Alaska, a pilot drops turkeys to rural homes for Thanksgiving, written by Mark Thiessen & Becky Bohrer, was originally published by Associated Press on November 28, 2024.
A.1 Finding Expert Annotators
The annotations for understanding the limitations of humans to detect AI-generated texts were done by 5 annotators who passed performance requirements based on the article expiriment of subsection 2.1. Only one annotator from the original experiment met these requirements. We recruited 10 more native English speakers on Upwork to take a 5 question sample of the original article task, questions which at most 2 out of 5 of the original annotators got correct. These 10 recruited annotators all had experience editing LLM generated content, frequently used LLMs, and had a professional background in editing or writing. Those who got at least 4 out of 5 correct (80%) given the rest of the 60 articles, and annotators who got at least 90% correct (54/60) were considered to be experts and recruited for the rest of the labeling rounds. 5 out of 10 annotators passed the 5 question trial; out of those 5, 4 passed the 60 question set.
Appendix B Dataset
In this section of the appendix we provide more details on our article corpus (B.1), how our AI articles were generated (B.2).
B.1 Article Corpus
Here we include more details about the articles collected for this study. Table 6 lists all publications of articles included in the corpus, [^22] All publications in the corpus were purchased by the researchers. with section distribution presented in Figure 7. Table 7 provides the statistics for articles by publication.
Publication Example Sections Pub. Date Range
Associated Press Science, Oddities, Animals May 15, 2024 - Dec 5, 2024
Discover Magazine Mind, The Sciences, Environment, Planet Earth July 10, 2024 - Nov 16, 2024
National Geographic Animals, Environment, Science, History & Culture, Travel Feb 8, 2023 - Nov 19th, 2024
New York Times US News, Science, Travel, Arts July 19, 2024 - Dec 6, 2024
Readers Digest Knowledge, Holidays March 16, 2023 - Nov 15, 2024
Scientific American Mind & Brain, Social Sciences, Technology May 3, 2024 - Nov 22, 2024, 2024
Smithsonian Magazine Smart News, Mind & Body, History, Innovation, Travel, Science Oct 7, 2022 - Dec 8, 2024
Wall Street Journal Science, Personal Technology, Workplace July 7, 2023 - Oct 29, 2024
Table 6: List of publications included in Human Detectors. The section is provided as listed as the section of the publication website where the article was published. All articles were taken from publications that wrote using American English.
Figure 7: Section distribution of articles across all trials.
Articles
Associated Press
(n=15)
Discover
(n=20)
National Geographic
(n=25)
New York Times
(n=20)
Readers Digest
(n=15)
Scientific American
(n=15)
Smithsonian Magazine
(n=25)
Wall Street Journal
(n=15)
Tokens (tiktoken)
Mean 860.0 949.5 1070.7 1009.1 865.6 874.2 922.0 927.5
St. Dev 240.7 229.9 203.2 138.4 171.9 347.1 210.7 202.5
Max 1236.0 1463.0 1497.0 1254.0 1318.0 1760.0 1454.0 1336.0
Min 442.0 561.0 710.0 645.0 634.0 455.0 532.0 608.0
Words (whitespace)
Mean 676.4 734.6 809.7 784.1 675.3 684.2 708.1 720.5
St. Dev 180.8 174.5 149.9 106.3 136.6 265.9 160.0 153.9
Max 1026.0 1117.0 1102.0 951.0 1008.0 1352.0 1095.0 1019.0
Min 342.0 415.0 536.0 513.0 501.0 362.0 388.0 491.0
Table 7: Number of tokens and words across articles by source.
B.2 AI Text Generation
We generate the articles by prompting the models with the articles title, subtitle, approximate length, publication, and section. For stories, we prompt models with the title of the reddit thread. All closed-source models were prompted using the provider’s API,  [^23] The estimated cost for generating all data with each model is as follows: GPT-4o \$1.85USD, o1-Pro \$3.81 USD, Claude-3.5-Sonnet \$0.51 USD. All models were prompted and articles with the prompt presented in Table 8. Instructions for rounding to an approximate length were included to ensure that articles on the same topic would be of similar length. Statistics on the distribution of lengths by trial are presented in Figure 8
Article Generation Prompt
You are given the following title and subtitle of a general article from the YOUR SECTION section and asked to write a corresponding article of around YOUR WORD COUNT words. Include quotations from relevant experts and make sure the article is concise and easily understandable to a lay audience.
Title: YOUR TITLE
Subtitle: YOUR SUBTITLE
Table 8: Prompt Template for Experiment 2 Paraphrasing
Figure 8: Word Count Distribution Per Experiment
Paraphrasing Attack
We use the sentence-level paraphrasing approach outlined in PostMark Chang et al. (2024), changing the exact language in the prompts slightly for our use case. This approach paraphrases the text on a sentence by sentence level. For the initial sentence, only that sentence is given to be paraphrased. For all following sentences, the portion of the text already paraphrased is passed to the LLM, intending to improve the overall flow of the paraphrased article. We use GPT-4o as the paraphraser model with a temperature set to 0. The prompt for the initial sentence can be found in Table 9 and the prompt for all following sentences can be found in Table 10.
Paraphrase Prompt (Initial Sentence Only)
Paraphrase the given sentence. Only return the paraphrased sentence in your response. Make it seem like a human wrote the article and that it is from the YOUR SECTION section of YOUR PUBLICATION.
Sentence to paraphrase: YOUR SENTENCE
Your paraphrase of the sentence:
Table 9: Prompt Template for Experiment 2 Paraphrasing
Paraphrase Prompt
Paraphrase the given sentence. Only return the paraphrased sentence in your response. Make it seem like a human wrote the article and that it is from the YOUR SECTION section of YOUR PUBLICATION.
Previous context: YOUR PARAPHRASED ARTICLE SO FAR
Sentence to paraphrase: YOUR SENTENCE
Your paraphrase of the sentence:
Table 10: Prompt Template for Experiment 2 Paraphrasing
Humanization Efforts:
To humanize the articles in §2.5, we tried many prompt-based approaches before settling on a prompting framework that was capable of evading the first author’s own AI-detection skills. Initially, we tried the following approaches that humanized already generated texts:
  1. 1.
    Generate, then Humanize: Firstly, we tried to replicate what we believe most students or novice LLM users would do to humanize their AI-generated text. We asked o1-Pro to first generate an article using the prompt template depicted in Table 8, then instructed the model to make it sound more human. Qualitative analysis from the first author found this method to result in an effect similar to paraphrasing.
  2. 2.
    Step-by-Step: To add a fine-grained approach to the humanization efforts, we next tried a step-by-step approach to humanization. We generated a base article, then iteratively prompted the model to humanize the article, focusing on a different element at each step. For example, we first asked it to make the article more creative, then alter the tone. Final steps included editing for grammar and replacing typical AI vocab. While this was better, the first author observed that the numerous calls to LLMs was adding more characteristics of AI, so we abonded this method.
  3. 3.
    Two-step Humanization: Next, we tried taking a 2-step approach to humanization. We gave the humanizing LLM the AI detection guide found in Table 11, and asked the LLM to make a list of everything identifiable as AI-generated in the article, and suggested edits. Then, giving the LLM the list of suggested edits, we prompted the LLM to edit the article to make it sound more human-written. Out of the humanization efforts, this yielded the best results.
Evader Details:
While approaches to humanization were improving, we found many of the largest identifiable traits of AI were due to the base generation. We decided to switch our approach from humanizing already existing AI-generated text to having an LLM generate the article in a humanized fashion all in one step. Our final prompt includes the set of instructions, examples of reference articles, detection guide, and the initial article prompt. The evader prompt template can be found in Table 13. The examples used were human-written and AI-generated articles from Experiments 1,2,3, and 4 from the same source. For example, when evading an article from New York Times, we provided all human-written New York Times articles and AI-generated articles based off those article titles.
Guide to Distinguishing AI-Generated Text from Human Writing Template
## Vocabulary / Word Choice Patterns
- Certain words crop up unusually frequently throughout AI-generated text compared to human writing
- Words like ‘delve’ and ‘tapestry’ are overused in AI-generated text but infrequently used in human writing ...
## Grammar
- Human writing generally less strictly adheres to English grammar rules and punctuation than AI-generated text
- AI-generated text uses a very formal writing style unless explicitly told not to ...
## Sentence Structure
- AI-generated sentences often follow the complex sentence structure, with multiple dependent and independent clauses, while human writing contains more of a mixture of simple, complex, and compound sentences
- example AI-generated sentence: “When it comes to celebrating Halloween, this holiday is a testament to the importance of empathy and community.” ...
## Formatting
- When AI makes lists, it typically uses the format of creating a bold header per bullet point, followed by a colon and then description of that list item.
- If a book title is referenced in a text, AI-generated text will always italicize the title, while human writing does not always follow this convention.
- Some pieces of text, especially articles and essays, contain headers and sub-headers. The headers written by AI are quite repetitive ...
## Tone
- The tone of AI-written text is flowery and formal, and its sentences are frequently structured as a reflective, onlooking statement, regardless of topic.
- AI tends to be inherently positive, attempting to emotionally uplift the reader, especially towards the conclusion.
- AI prioritizes efficiency, sometimes sacrificing clarity or depth in its messaging ...
## Introductions
- AI-written introductions often contain a strong scene-opener with a description of a specific time or place, such as "On a drab November morning..." or "On December 8, 1660, a London audience gathered ..." ...
## Conclusions
- AI-generated text always ends with a neat conclusion, instead of just ending the article naturally.
- AI-generated conclusions are often overly long and summarize everything that has already been written in an article ...
## Content
- Unless specifically prompted, AI will avoid controversial topics at all costs. - AI will avoid any type of swear word, including mild ones like ‘darn’, or any other offensive vocabulary ...
## Contextual Accuracy and Factuality
- Human writing in the domain of non-fiction is factually accurate and contains many specific factual claims.
- In human writing, people, places, brands, and other named objects can be verified or are highly plausible ...
## Creativity & Originality
- AI-generated text is much less creative than that of humans, lacking originality and sticking to an ‘obvious’ way to answer a prompt.
- Humans incorporate twists, unexpected insights, and twists that AI hasn’t seemed to master quite yet ...
Table 11: A truncated version of the AI Text Detection Guide. The full guide is located at https://github.com/jenna-russell/human_detectors.
AI Vocabulary Included in Detection Guide
Nouns: aspect, challenges, climate, community, component, development, dreams, environment, exploration, grand scheme, health, hidden, importance, landscape, life, manifold, multifaceted, nuance, possibilities, professional, quest, realm, revolution, roadmap, role, significance, tapestry, testament, toolkit, whimsy
Verbs: capturing, change, consider, delve/dive into, elevate, embrace, empower, enact, enhance, engage, ensure, evoking, evolving, explore, fostering, guiding, harness, highlights, improve, integrate, intricate, jeopardizing, journey, navigating, navigate, notes, offering, partaking, resonate, revolutionize, shape, seamlessly, support, tailor, transcend, underscores, understanding
Adjectives: authentic, complex, comprehensive, crafted, creative, critical, crucial, curated, deeper, diverse, elegant, essential, groundbreaking, key, meaningful, paramount, pivotal, powerful, profound, quirky, robust, seamless, significant, straightforward, structured, sustainable, transformative, valuable, vast, vibrant, vivid, whimsical
Adverbs: additionally, aptly, creatively, moreover, successfully
Phrases: as we [verb] the topic, cautionary tale, connect with, has shaped the, in a world of/where, in conclusion, in summary, it’s crucial to, it’s important to note, it’s not about ___ it’s about ___, manage topic issues/problems, not only ___ but also, packs a punch/brings a punch, paving the way, personal growth, quality of life, remember that, simple yet ___, step-by-step, such as, the effects of, the rise of, their understanding of, they identified patterns, to form the, to mitigate the risk, weaving, when it comes to topic
Table 12: All ‘AI’ Vocabulary our expert annotators noted, as listed in the Detector Guide. See the full detection guide prompt in Table 11.
Evader Template
You are a journalist writing an article for . Your goal is to write an article given a title and subtitle, but make it look like it was written by a human. To assist you, we provide the following detection guide to determine if something is written by a human or by AI.
AI Detection Guide
You are given some examples of human and AI written articles to help you.
<start examples>
Example Articles
<end examples>
You are also given the following guidelines to make your article seem like a human wrote it.
<start of guidelines>
- Follow the detection guideline to ensure writing does not sound AI-generated.
- Do not use any of the "Overused AI Words/Phrases", "Overused AI Metaphors", "AI Grammar Patterns", AI sentence structures, AI tones, common AI names, or AI content described in the guide.
- Include words/phrases/grammar patterns/tone/content that are consistent with human writing as described in the guide.
- Use human-written examples and explanations as references for formatting and content.
- Use the example explanations to see what humans identify as ’AI’ or ’Human’
- Avoid formulaic introductions that elaborate on scene-setting or time, immediately quote experts, or bring in too much historical context.
- Avoid formatting the first setting by mentioning the setting, then a comma, then some detail or context. Avoid mentioning atmospheric details or time (of day, month, year, etc ...) in the first sentence as well.
- Avoid generic, vague, and forward-looking conclusions, and do not summarize the article in the conclusion to make articles sound more human-like. Concluding sentences don’t have to wrap everything up nicely.
- Avoid using generic statements to start paragraphs, such as ’For now’ or ’In the end’
- Add references to darker topics when appropriate.
- Add specific references to places, people, brands, items, facts, and metrics when appropriate.
- Always opt for specifics over broad details. For example, mention correct statistics instead of broad claims and mention company names instead of saying the type of company.
- Make sure that none of the claims made in your article are factually incorrect or implausible.
- Use language aligned with the publication you are representing as a journalist
- If your article would benefit from being organized into sections (e.g., scientific or historical content), please use section headers to do so. There are 1/3 odds an article has headers.
<end of guidelines>
Now, by following the guidelines above, examples, and the information in the detection guide, please write an article without any AI signatures that also includes signatures of human writing. No reader should doubt that your article could have been published by the given source. Readers with access to the information in the detection guide should not be able to detect the article was written by AI.
Article Generation Prompt
Article:
Table 13: Prompt used for evader. The first insert into the prompt is filled by Table 11. The second insert is filled by examples of human and machine-generated articles.
New Telescope Could Potentially Identify Planet X
Are there hidden planets in our solar system? New technologies, like the powerful Rubin Observatory, brings us closer to answers.
For decades, astronomers have speculated that there might be another large planet lurking in the outer reaches of our solar system—an elusive “Planet X.” Often referred to as “Planet Nine” by modern researchers, this hypothetical world is believed to be so distant that it has escaped detection by existing telescopes, leaving only subtle gravitational clues in its wake. Now, a new generation of telescopes, particularly the upcoming Vera C. Rubin Observatory, promises a more powerful set of eyes on the night sky. With these tools, scientists hope to either pinpoint this mysterious planet’s location or finally put the idea to rest.
Why the Search Matters
The concept of a hidden planet is not new. In the 19th century, the discovery of Neptune followed suspicions that Uranus’s odd orbital path hinted at an unseen gravitational pull. Today, attention has turned to the distant suburbs of our solar system, where far-off objects in the Kuiper Belt—frozen remnants of planetary formation—appear to cluster in strange ways. Some astronomers argue that only a large, unseen planet’s gravity could explain these odd orbits.
“The evidence is subtle, but it’s there,” said Dr. Konstantin Batygin, a planetary astrophysicist at the California Institute of Technology (Caltech) who, along with colleague Dr. Mike Brown, first proposed the existence of a Planet Nine in 2016. “We’re seeing several distant Kuiper Belt objects all tilted and clustered in a peculiar manner, and a ninth planet several times Earth’s mass, orbiting far beyond Neptune, could be the simplest explanation.”
Not everyone is convinced. Skeptics point to the small number of known distant objects and argue that the clustering could be a statistical fluke. Others think unseen observational biases might make it look like these objects are behaving strangely. Either way, the mystery remains unsolved—at least for now.
Enter the Vera C. Rubin Observatory
Scheduled to begin its full operations in the near future, the Vera C. Rubin Observatory (formerly known as the Large Synoptic Survey Telescope, or LSST) in Chile represents a leap forward in astronomical capability. With a massive 8.4-meter mirror and a state-of-the-art digital camera, it will repeatedly scan the entire southern sky over the course of a decade, providing a dynamic, time-lapse-like portrait of celestial motions.
“The Rubin Observatory is really a game-changer,” said Dr. Meg Schwamb, an astronomer at Queen’s University Belfast who studies the outer solar system. “Instead of looking at a small patch of sky, we’ll be looking at pretty much everything visible from Chile, over and over again. This repeated coverage means we can detect faint, distant objects that move slowly across the sky—exactly the kind of signature we’d expect from a far-flung planet.”
By systematically imaging the sky every few nights, the Rubin Observatory’s Legacy Survey of Space and Time (LSST) will reveal thousands—or even tens of thousands—of new objects in the outer solar system. Among them could be Planet Nine, if it exists. Even if the telescope doesn’t directly spot the planet, it might detect more distant objects whose orbits can be mapped with unprecedented precision, allowing astronomers to figure out once and for all if a hidden giant is lurking out there.
Technological Edge
One reason Planet Nine (or any Planet X) has been so hard to pin down is that if it exists, it’s incredibly faint and slow-moving, possibly hundreds of times farther from the Sun than Earth is. Traditional telescopes rely on painstaking surveys that cover only small portions of the sky at a time. In contrast, Rubin’s wide field of view—about 40 times the size of the full Moon—means it will cover the visible sky every few days.
“This isn’t just another telescope—it’s a new way of doing astronomy,” said Dr. Lynne Jones, an LSST researcher at the University of Washington. “We’re moving from static snapshots to continuous movies. If there’s something out there, no matter how faint, as long as it’s moving, we have a good chance of picking it up over time.”
Ruling In or Ruling Out
If the Rubin Observatory does find Planet Nine, the discovery would reshape our understanding of the solar system’s architecture. A distant giant could be a leftover core of a gas giant that got tossed out in the early days of planetary formation, or it might have formed in situ, far from the Sun’s warmth. Such a find could help scientists piece together the chaotic early era when planets were jostling for position and smaller bodies were flung into distant orbits.
On the other hand, if Rubin’s comprehensive survey concludes after several years without any sign of Planet Nine—or if newly discovered distant objects don’t line up as predicted by the planet hypothesis—then the case for a hidden giant will weaken. Instead, astronomers might refine models of how solar system objects distribute themselves naturally. Understanding these patterns could still teach us valuable lessons about how gravity and cosmic debris shape the outskirts of our celestial neighborhood.
Beyond Planet X
The quest for Planet Nine is just one highlight of what the Rubin Observatory offers. It will also track near-Earth asteroids that could pose future hazards, study dark matter and dark energy by observing the distribution of galaxies, and contribute to countless other fields of astronomy.
The quest for Planet Nine is just one highlight of what the Rubin Observatory offers. It will also track near-Earth asteroids that could pose future hazards, study dark matter and dark energy by observing the distribution of galaxies, and contribute to countless other fields of astronomy.
Still, for many sky-watchers, the idea of a hidden planet holds a particular allure. The notion that our solar system might still hold major surprises underscores how much we have yet to learn about our own cosmic backyard.
“As much as astronomy has advanced, we’re still explorers,” said Batygin. “The Rubin Observatory gives us the tools to either find this planet or lay the mystery to rest. In the next decade, I expect we’ll know a lot more about what’s really out there—and what isn’t.”
Until then, the search continues, driven by powerful new telescopes and the human desire to uncover the unseen.
Table 14: The one article the majority of annotators did not detect correctly, generated from o1-Pro as part of Experiment 4 (see §2.4). "New Telescope Could Potentially Identify Planet X" was originally written by Emilie Le Beau Lucchesi in Discover Magazine on Nov. 6th, 2024
Appendix C Pilot Study
In this section, we provide details on the pilot study, where five annotators with varying frequencies of familiarity with LLMs detect if texts are human-written or AI-generated. The pilot study follows the same annotation process described in §2. In the study, annotators completed two rounds of testing, one using the same articles as identified in §2.1, another using stories from the subreddit r/WritingPrompts.
Story Corpus:
Here we include more details about the stories collected for this study. We collect 30 stories from r/WritingPrompts. We generate corresponding AI-generated stories with the prompt in Table 15.
Story Generation Prompt
Please write a story around WORD COUNT words responding to the following prompt: STORY PROMPT.
Table 15: Prompt Template for Story Generation, where STORY PROMPT is the writing prompt from r/WritingPrompts that the human-written story was written about and WORD COUNT is the length of the story to generate.
Story Results :
Overall, nonexperts faired better on stories than articles, with an average TPR of 69.3% (including performance of the expert annotator). The nonexperts had a TPR of 62.5% not including our expert annotator, as shown in Table 16. We found that many of the differences spotted were superficial due to the very informal nature of reddit, compared to our prompt which did not instruct the LLM to generate stories of that nature. In reddit, people use acronyms, for example shortening ‘you’ to ‘u’, type in ALL CAPS, or generally write in a very R-rated manner. While these are real differences between AI-generated and human-written texts, we wanted to pursue clues humans could use in settings where a person was genuinely attempting to pass of AI-generated writing as human-writtten, such as students writing an essay or someone publishing a news article. Future work may explore human detection abilities on fictional work, where the human-written references are more edited than the reddit thread stories.
Annotator TPR (%) FPR (%)
Nonexpert Annotator 1 73.3 76.7
Nonexpert Annotator 2 43.3 80.0
Nonexpert Annotator 3 66.7 80.0
Expert Annotator 1 96.7 93.3
Nonexpert Annotator 4 66.7 76.7
Average 69.3 81.3
Table 16: Story performance of initial 5 annotators, which included 4 nonexpert annotators and expert annotator # 1, who was used in all article experiments.
Appendix D Comment Analysis
In this section, we outline the framework for analyzing expert explanations and provide additional analysis of explanations.
D.1 Categorization of Comments
To categorize explanation comments, we first define the explanation categories found in Table 4. The first two authors individually annotated a sample of 25 expert explanations. The sample is a stratified sample of one comment per expert per each of the five experiments. Authors then came to an agreement on final human labels, refining categories as needed. We then prompt GPT-4o-2024-11-20 with the prompt found in Table 17 to categorize the sample explanations. Once the prompt was able to classify the majority of the sample explanations in alignment with the authors labels, we prompt GPT-4o-2024-11-20 to categories the explanations of all 1500 explanations. Each explanation can contain labels for multiple categories, since most explanations touch on multiple reasons of why a text is AI-generated. The total cost of classifying comments using GPT-4o was \$6.02 USD.
Comment Categorization Prompt
We hired an annotator to determine whether an article is AI-generated or human-written. Alongside their “machine-generated” or “human-generated” label, they provided an **explanation** detailing the specific clues that led them to their decision.
Your task is to **review** the annotator’s explanation and **categorize** each clue they mention. For each clue, identify:
1. The **category** it falls under (e.g., Vocabulary, Grammar, etc.).
2. The **label** (AI-generated or Human-written) the annotator associates with that clue.
3. The exact **quote** from the annotator’s explanation that shows this clue.
Below is a list of categories and **example** indicators the annotator might reference. **Note:** These examples do **not** cover all possibilities. If a comment fits two categories, choose the best one.
{DEFINITIONS_OF_ALL_CATEGORIES}
Instructions:
- Read the annotator’s explanation.
- Identify any category that applies.
- Indicate whether the annotator says it points to AI-generated or human-written.
- Provide the exact quote from the annotator’s explanation that led you to this conclusion.
**If the annotator cites both AI and human clues within the same category**, please create **two separate entries** (one for AI-generated, one for human-written).
**If no category is applicable**, categorize it under “other.”
**Answer Format** (use this structure for each separate category/label pair):
<category>YOUR CATEGORY HERE</category>
<label>YOUR LABEL HERE (AI-generated or Human-written)</label>
<quote>RELEVANT QUOTE FROM EXPLANATION</quote>
<category>YOUR CATEGORY HERE</category>
<label>YOUR LABEL HERE (AI-generated or Human-written)</label>
<quote>RELEVANT QUOTE FROM EXPLANATION</quote>
(Repeat as needed, using a new block for each category/label pair.)
Example:
<category>Vocabulary</category>
<label>AI-generated</label>
<quote>"The article mentioned the word ‘crucial’ and used a lot of unusual synonyms."</quote>
The annotators explanation is as follows:
<explanation>EXPLANATION</explanation>
Table 17: Truncated prompt used for comment analysis. The first insert is filled by a list of definitions of the categories found in Table 4 and the second insert is filled by an annotator explanation. A full version of the guide can be found at https://github.com/jenna-russell/human_detectors.
D.2 Individual Experts Commentary
Each expert had clues they favored using throughout all experiments. Annotator 1,whose category mention frequencies can be found in Figure 9, Figure 10 depicts Annotator 2 comments, Figure 11 shows Annotator 3 comments, Figure 12 shows Annotator 4 comments and Figure 13 has commentary frequencies from Annotator 5. The individual heatmaps highlight the range of clues used by annotators, who individually had no reference of clues other experts used.
Figure 9: Annotator 1 Frequency of Categories Mentioned in Explanations
Figure 10: Annotator 2 Frequency of Categories Mentioned in Explanations
Figure 11: Annotator 3 Frequency of Categories Mentioned in Explanations
Figure 12: Annotator 4 Frequency of Categories Mentioned in Explanations
Figure 13: Annotator 5 Frequency of Categories Mentioned in Explanations
Appendix E Automatic Detection
E.1 Explainable Detection
In this section we detail the prompt based detector used for experiments in §4.
Implementation Details
The prompt-based detector is set up to mimic how humans think about AI-generated text detection rather rather than how current automatic detectors do.
  • Zero-Shot: To find a baseline performance of detector models, we prompt the model using the template in Table 18 to return if the candidate text is Human-written or AI-generated.
  • Zero-Shot + COT: We prompt the model using the template in Table 19 to return if the candidate text is Human-written or AI-generated and an explanation of why they text is human-written. This ablation was conducted to understand the effect explanations may have on over LLM detection performance.
  • Zero-Shot + Guide: In this experiment, we prompt the model using the template in Table 20 to return if the candidate text is Human-written or AI-generated. This experiment serves to understand the effect including the guide in the prompt has on LLM performance.
  • Zero-Shot + COT + Guide: In this experiment, we prompt the model using the template in Table 21 to return if the candidate text is Human-written or AI-generated and an explanation of why they text is human-written. THis set up is set up to fully mimic how a human detects, thinking through clues and providing explanations of what makes them think a text is either human-written or AI-generated.
We observe that expert annotators perform better when using a majority vote strategy. Future work in prompt-based detectors could employ majority voting as a strategy to increase detection accuracy and avoid false positives.
Zero Shot Detector Template
You are given a candidate text and an AI detection guide. Your task is to carefully read the candidate text and determine whether it was either written by a human or generated by AI.
Answer HUMAN-WRITTEN if the candidate text was likely written by a human.
Answer AI-GENERATED if the candidate text was likely generated by an AI.
<start of candidate text> Candidate Text </end of candidate text>
<question>Is the above candidate text HUMAN-WRITTEN or AI-GENERATED? </question>
Provide your final answer in this format.
<answer> YOUR ANSWER </answer>
Table 18: Prompt Template for the zero-shot detector set up
Zero-shot + CoT template
You are given a candidate text. Your task is to carefully read the candidate text and determine whether it was either written by a human or generated by AI.
Answer HUMAN-WRITTEN if the candidate text was likely written by a human.
Answer AI-GENERATED if the candidate text was likely generated by an AI.
<start of candidate text> Candidate Text </end of candidate text>
<question>Is the above candidate text HUMAN-WRITTEN or AI-GENERATED? </question>
First, concisely describe the features of the candidate text that exemplify either AI or human writing. Then, provide your final answer.
<description> YOUR DESCRIPTION </description>
<answer> YOUR ANSWER </answer>
Table 19: Prompt Template for the Zero-shot + CoT detector setup
Zero-shot + Guide Template
You are given a candidate text and an AI detection guide. Your task is to carefully read the candidate text and determine whether it was either written by a human or generated by AI.
Answer HUMAN-WRITTEN if the candidate text was likely written by a human based on the information in the provided guide.
Answer AI-GENERATED if the candidate text was likely generated by an AI based on the information in the provided guide.
DETECTION GUIDE
<start of candidate text> Candidate Text </end of candidate text>
<question>Based on the detection guide, is the above candidate text HUMAN-WRITTEN or AI-GENERATED? </question>
<answer> YOUR ANSWER </answer>
Table 20: Prompt Template for the Zero-shot + Guide detector setup
Zero-shot + CoT + Guide Template
You are given a candidate text and an AI detection guide. Your task is to carefully read the candidate text and determine whether it was either written by a human or generated by AI.
Answer HUMAN-WRITTEN if the candidate text was likely written by a human based on the information in the provided guide.
Answer AI-GENERATED if the candidate text was likely generated by an AI based on the information in the provided guide.
DETECTION GUIDE
<start of candidate text> Candidate Text </end of candidate text>
<question>Based on the detection guide, is the above candidate text HUMAN-WRITTEN or AI-GENERATED? </question>
First, use the provided guide to concisely describe the features of the candidate text that exemplify either AI or human writing. Then, provide your final answer.
<description> YOUR DESCRIPTION </description>
<answer> YOUR ANSWER </answer>
Table 21: Prompt Template for the Zero-shot + CoT + Guide detector setup
E.2 Results
Thresholding for Automatic Detectors:
Some automatic detectors benchmarked in Table 3 and Table 22 did not provide suggested thresholds for model usage. In these scenarios, we test models on a held-out test set 40 human-written articles, finding a threshold for a 5% FPR. The respective thresholds used were 0.6051510572 for RADAR, 0.96 for Fast-DetectGPT, and 0.8963184953 for e5-lora.
Detector Model
We initially test GPT-4o-2024-08-06, GPT-4o-2024-11-20, o1, and Claude-3.5-Sonnet as potential detector models. Based on early results, we continue with ablations of ø1 due to its low FPR and GPT-4o-2024-11-20 due to having the highest TPR. The cost of all o1 detection experiments on the 300 texts is \$100.60 USD and \$25.24 USD for GPT-4o-2024-11-20 detection experiments.
Additional Results:
While testing both GPT-4o-2024-08-06 and GPT-4o-2024-11-20 as detectors, we find extremely varied performance, with GPT-4o-2024-11-20 performing much better, particularly at detecting paraphrased and o1 articles. Claude-3.5-Sonnet failed to detect content generated by itself and o1Ẇe also tested e5-lora, one of the top performing models on the Raid Dugan et al. (2024) benchmark, which we found was unable to correctly classify any AI-generated articles at a FPR of 5%. See all additional results in Table 22.
[…]